OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. r = 6

OpenStudy (lgbasallote):

\[Area = 3.14 \times (radius)^2\]

OpenStudy (anonymous):

It would be 113.04 right?

OpenStudy (lgbasallote):

that means \[Area = 3.14 \times (6)^2\] \[\Rightarrow Area = 3.14 \times 36\] got that??

OpenStudy (anonymous):

yeah!

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

113.04 is right ^_^

OpenStudy (anonymous):

thank you!!